Find the compound amount and the amount of interest earned by the following deposit. ​$2,000 at 5.41​% compounded continuously for 6 years.

What is the compound​ amount? ​$ ​(Round to the nearest​ cent.)

What is the interest earned?
Answer #1

Future Value = Present Value * e^(Rate per year*no. of years)

= 2000*e^(.0541*6)

= 2000*e^0.3246

= 2000*1.38347714443

2766.95

interest earned = Future Value - Present Value

= 2766.95-2000

766.95